find rate of change of f(x)=x^2+3x+2
X=0 X=4

Answer by nerdybill(7384)

find rate of change of f(x)=x^2+3x+2
X=0 X=4
.
first derivative:
f'(x)=2x+3
f'(0)=2(0)+3 = 0+3 = 3
f'(4)=2(4)+3 = 8+3 = 11